# Numerical approximation of McKean-Vlasov SDEs via stochastic gradient descent

### v2

@@ 3       ${\bf a}_{m+1}:={\bf a}_{m}-\eta_{m}{\bf v}_{m+1}$;
The function $v$ is defined component-wise by (3.1). Note that the $\mathbb{R}^{(n+1)K}$-valued stochastic processes $({\bf a}_{m})_{m\in\mathbb{N}}$ and $({\bf v}_{m})_{m\in\mathbb{N}}$ are recursively defined, as in Algorithm 1, and are adapted to the filtration (¯$(\bar{\mathcal{F}}_{m})_{m\in\mathbb{N}}$, which then represents the history of the algorithm.
@@
Input : Maximum number of iterations $m_{\max}$; Mini-batch sample size $M$
@@ 3.3 Convergence of Algorithm 1
where we set ¯$\bar{\mathcal{F}}_{0}:=\{\bar{\Omega},\emptyset\}$. Furthermore, the random variables
@@
Differentiability of $G$ follows from Proposition 3.4. By Proposition 3.3 and Proposition 3.4, the hypothesis of [7, Proposition 3] and [35, Theorem 1] are fulfilled. In particular, Theorem 1 in [35] (see the remark on p.5 there) implies that $\bar{\mathbb{P}}$-almost surely either ${{\bf a}_{m}}$ converges in $\mathbb{R}^{(n+1)K}$ or $|{{\bf a}_{m}}|\to+\infty$. On the other hand, by Proposition 3 in [7], either $G({\bf a}_{m})\to-\infty$ or $G({\bf a}_{m})$ converges to a finite value and $\nabla G({\bf a}_{m})\to 0$, $\bar{\mathbb{P}}$-almost surely. Since
@@ 4 Numerical study
We conduct several numerical experiments to study the numerical performance of Algorithm 2 for different values of the mini-batch sample size $M$. Precisely, we test it for the following models:
@@
    $\varepsilon_{m}:=\frac{\|\mathscr{L}{\bf{a}}_{m}-{\bar{\gamma}}^{\text{MC}}\|_{L^{2}_{T}}}{\|{\bar{\gamma}}^{\text{MC}}\|_{L^{2}_{T}}},$ (58)
    with $\mathbf{a}_{m}$ being the $m$-th iterate of Algorithm 2 and the $L^{2}_{T}$ norm over the interval $[0,T]$ is as defined in Table 1. The algorithm is always halted when the error $\varepsilon_{m}$ as computed above is below 1%.
@@ 4.1 Kuramoto-Shinomoto-Sakaguchi MV-SDE
Consider the following MV-SDE, which is related to the famous Kuramoto-Shinomoto-Sakaguchi model (see [20, Section 5.3.2]):
@@
We test different choices of algorithm hyperparameters: $r_{0}=1,5,10$, $\rho=0.6,0.7,0.8,0.9$, $n=3,4,5,6{,7,8}$ and $M=1,10,100,1000,10000$. The benchmark ${\bar{\gamma}}^{\text{MC}}$ in (57), based on the Monte Carlo particle system approximation, is obtained with $N_{0}=10^{6}$ particles. For each combination of hyperparameters, we perform 1000 independent runs of the algorithm, and record the average number of iterations of the SGD algorithm required to achieve the relative error accuracy $\varepsilon_{m}<1\%$.
@@
In Figures 2, 3, 4 we plot the output curves $\mathscr{L}{\bf a}_{m}=(\mathscr{L}{\bf a}_{m}^{1},\mathscr{L}{\bf a}_{m}^{2})$ of the SGD algorithm versus the benchmark curves ${\bar{\gamma}}^{\text{MC}}=({\bar{\gamma}}^{\text{MC}}_{1},{\bar{\gamma}}^{\text{MC}}_{2}),$ for all values of the polynomial degree $n=3,4,...,8$ and for time horizons $T=0.5,1,2,$ respectively. Note that the approximation does not deteriorate as the time horizon $T$ increases. In particular, the slight loss of precision that can be observed towards the right end of the time interval can be explained by the choice of polynomial approximations. For instance, the gap between $(\mathscr{L}{\bf a}_{m})^{1}$ and ${\bar{\gamma}}^{\text{MC}}_{1}$, which can be seen in Figure 2 in the region $t\in[0.4,0.5]$, does not appear in the analogous plot of Figure 3 (or Figure 4) after the same number of steps. This shows that the error does not propagate over time: this is mainly due to the fact that ours is a full history-type minimization algorithm, in that it approximates the minimum in (11) over the set of the continuous curves on $[0,T]$.
@@ 4.2 Polynomial drift MV-SDE
In Figure 5 we plot the output curves $\mathscr{L}{\bf a}_{m}=(\mathscr{L}{\bf a}_{m}^{1},\mathscr{L}{\bf a}_{m}^{2})$ of the SGD algorithm versus the benchmark curves ${\bar{\gamma}}^{\text{MC}}=({\bar{\gamma}}^{\text{MC}}_{1},{\bar{\gamma}}^{\text{MC}}_{2})$ for all time horizons $T=0.1$, $T=0.5$ and $T=1$. In the plot for $T=0.5$, notice the loss of accuracy in our approximation $\mathscr{L}{\bf a}_{m}^{2}$ of $\bar{\gamma}_{2}(t)=\mathbb{E}[X_{t}]$ over the period $[0.45,0.50].$ This is akin to Runge’s phenomenon in polynomial interpolation over compact sets, and is possibly due to the norm used in the computation of $\varepsilon_{m}$ in (58). One possibility of improvement is to first employ the alternative objective function presented in Remark 2.14, i.e., using a weight function to force the SGD algorithm to be more accurate towards the end of the interval. Additionally, one can also use a similar weight function to compute a weighted $L^{2}_{T}$ norm for error $\varepsilon_{m}$ in the stopping criterion.
@@ 4.3 Convolution-type MV-SDE
[…]
@@
In Figure 6 we plot the approximate density $\tilde{w}^{(K),\text{MC}}_{T}(x)$ for $K=3,5,10,20$, with $x$ ranging on the interval $[-3,4]$ From this we conclude that the value $K=10$ is high enough in order for $\tilde{w}^{(K)}$ to accurately approximate ${w}_{T}$ over the interval $[-3,4]$. In Figure 7 we test the accuracy of the SGD algorithm by plotting $\tilde{w}^{(K),\text{SGD}}_{T}(x)$ versus $\tilde{w}^{(K),\text{MC}}_{T}(x)$, with $K=10$, over the same interval. The parameters of the SGD algorithm were chosen as $n=3$, $M=100$, $r_{0}=5$, $\rho=0.9$, and the algorithm halted after $m=172$ iterations ($35$ seconds of computation time) when the relative error reached $\varepsilon_{m}<1\%$.
Figure 6: Monte Carlo method densities $\tilde{w}^{(K),\text{MC}}_{T}(x)$ over the interval $[-3,4]$, for $T=1$ and $K=3,5,10,20$. The benchmark vector ${\bar{\gamma}}^{\text{MC}}(T)$ was computed with $N_{0}=10^{7}$ particles. Here $X_{0}\sim\mathcal{N}_{(0,1)}$ and $\sigma=0.1$. In the model (63) we had $X_{0}\sim\mathcal{N}_{(0,1)}$ and $\sigma=0.1$.
Figure 7: Comparison of $\tilde{w}^{(K),\text{SGD}}_{T}(x)$ and $\tilde{w}^{(K),\text{MC}}_{T}(x)$ over the interval $[-3,4]$, for $T=1$ and $K=10$. The benchmark vector ${\bar{\gamma}}^{\text{MC}}(T)$ was computed with $N_{0}=10^{7}$ particles. The SGD output $\mathscr{L}{\bf a}_{m}(T)$ was obtained after $m=172$ iterations ($35$ seconds of computation time), with parameters $n=3$, $M=100$, $r_{0}=5$, $\rho=0.9$. In the model (63) we had $X_{0}\sim\mathcal{N}_{(0,1)}$ and $\sigma=0.1$.
@@ Appendix A SDE stability estimates
Furthermore, for a given $\kappa>0$ and for any $\rho$ and $\hat{\rho}$ with $\|\rho\|_{\mathcal{S}^{\infty}_{T}},\|\hat{\rho}\|_{\mathcal{S}^{\infty}_{T}}\leq\kappa$ we have:
@@
    $\lVert\mathcal{Y}^{\rho}-\mathcal{Y}^{\hat{\rho}}\rVert_{\mathcal{S}^{2p}_{[0,T]}}\leq C_{3}\,\|\rho-\hat{\rho}\|_{\mathbb{L}^{2p}_{[0,T]}}\leq C_{3}\,T^{\frac{1}{2p}}\|\rho-\hat{\rho}\|_{\mathcal{S}^{\infty}_{T}},$ (76)
@@
The well-posedness of $\mathcal{Y}^{\rho}$ and part (a) of the result are particular cases of [30, Theorem 2.5.7 (p.77)] and [30, Theorem 2.5.10 (p.85)], respectively.
Parts (b) and (c) can be readily obtained by combining (73) with [30, Theorem 2.5.9 (p.83)]. Precisely, from [30, Theorem 2.5.9], there exists $C>0$, only dependent on $\kappa,p,T,\Lambda$ and $K$, such that
@@ B.1 Proof of the time-regularity of $\bar{\gamma}$
Let $X$ be the unique solution to (2), which exists due to Proposition 2.4. A straightforward application of Itô’s formula shows that its time-marginal laws $\mu_{t}(\text{d}x)$ solve the Fokker-Plank equation in distributional form over $[0,T]\times\mathbb{R}^{d}$, which is
@@ B.2 Auxiliary stability and moment estimates results
$+\big{|}\nabla{{\bf h}_{j}}\big{(}\mathscr{L}a(t)\big{)}\beta(t,Z^{a}_{t})-\nabla{{\bf h}_{j}}\big{(}\mathscr{L}a^{\prime}(t)\big{)}\beta(t,Z^{a^{\prime}}_{t})\big{|}\Big{)}$ (96)
@@
Since ${\bf h},\alpha(t,\cdot),\beta(t,\cdot)$ and their derivatives up to order $2$ are bounded, uniformly w.r.t. $t\in[0,T]$, [31, Theorem 2.3.1 (p.218)] yields
[…]
@@
$\mathbb{E}\big{[}\varphi(Z^{a}_{t})\big{]}-\mathbb{E}\big{[}\varphi(Z^{0}_{t})\big{]}=\mathbb{E}\big{[}\varphi(Z^{a}_{t})-\varphi(Z^{0}_{t})\big{]}$ $=\mathbb{E}\bigg{[}\int_{0}^{a}\partial_{\hat{a}}\big{\{}\varphi(Z^{\hat{a}}_{t})\big{\}}\text{d}\hat{a}\bigg{]}$ (106)
@@ B.3 Proof of the main auxiliary propositions
Recall the heuristics in (45) and (3.1). For any independent copies $(\xi,W)$ and $(\tilde{\xi},\tilde{W})$, since ${\bf h},{\bf h}^{\prime},\varphi,\varphi^{\prime}$ are bounded, we observe that
$\mathbb{E}\big{[}\big{|}\varphi\big{(}Z^{a}_{t}(\bar{\xi},\bar{W})\big{)}-{\bf h}\big{(}(\mathscr{L}a)(t)\big{)}\big{|}^{2}\big{]}\times\mathbb{E}\big{[}\big{|}\varphi^{\prime}\big{(}Z^{a}_{t}(\tilde{\xi},\tilde{W})\big{)}Y^{a}_{t}(\tilde{\xi},\tilde{W})\big{)}-$ $\partial_{a}{\bf h}\big{(}(\mathscr{L}a)(t)\big{)}\big{|}^{2}\big{]}$
@@
$\mathbb{E}_{\bar{\mathbb{P}}}\big{[}{\bf v}_{m+1}-\nabla_{a}G({\bf a}_{m})|\bar{\mathcal{F}}_{m}\big{]}$ (114)
$=\mathbb{E}_{\bar{\mathbb{P}}}\Big{[}v({\bf a}_{m};\xi_{m+1},W_{m+1};\tilde{\xi}_{m+1},\tilde{W}_{m+1})-\mathbb{E}_{\bar{\mathbb{P}}}\big{[}v(a;\xi_{m+1},W_{m+1};\tilde{\xi}_{m+1},\tilde{W}_{m+1})\big{]}_{a={\bf a}_{m}}\Big{|}\bar{\mathcal{F}}_{m}\Big{]}=0,$ (115)

### v3

@@ 4 ${\bf a}_{m+1}:={\bf a}_{m}-\eta_{m}{\bf v}_{m+1}$;
The function $v$ is defined component-wise by (3.1). Note that the $\mathbb{R}^{(n+1)K}$-valued stochastic processes $({\bf a}_{m})_{m\in\mathbb{N}}$ and $({\bf v}_{m})_{m\in\mathbb{N}}$ are recursively defined, as in Algorithm 1, and are adapted to the filtration $(\bar{\mathcal{F}}_{m})_{m\in\mathbb{N}}$, which then represents the history of the algorithm.
@@
Input : Maximum number of iterations $m_{\max}$; learning rates $(\eta_{m})_{m\in\mathbb{N}}$; Mini-batch sample size $M$
@@ 3.3 Convergence of Algorithm 1
where we set $\bar{\mathcal{F}}_{0}:=\{\bar{\Omega},\emptyset\}$. Furthermore, the random variables
@@
Differentiability of $G$ follows from Proposition 3.4. By Proposition 3.3 and Proposition 3.4, the hypothesis of [6, Proposition 3] and [41, Theorem 1] are fulfilled. In particular, Theorem 1 in [41] (see the remark on p.5 there) implies that $\bar{\mathbb{P}}$-almost surely either ${\bf a}_{m}$} converges in $\mathbb{R}^{(n+1)K}$ or $|{{\bf a}_{m}}|\to+\infty$. On the other hand, by Proposition 3 in [6], either $G({\bf a}_{m})\to$-$\infty$ or $G({\bf a}_{m})$ converges to a finite value and $\nabla G({\bf a}_{m})\to 0$, $\bar{\mathbb{P}}$-almost surely. Since
@@ 4 Numerical study
We conduct several numerical experiments to study the numerical performance of Algorithm 2 for different values of the mini-batch sample size MM [^2] The code is available at https://github.com/amatoandre/Numerical-approximation-of-MKV-SDEs-via-SGD.git. Precisely, we test it for the following models:
@@
    $\varepsilon_{m}:=\frac{\|\mathscr{L}{\bf{a}}_{m}-{\bar{\gamma}}^{\text{MC}}\|_{L^{2}([0,T])}}{\|{\bar{\gamma}}^{\text{MC}}\|_{L^{2}([0,T])}},$ (58)
    with $\mathbf{a}_{m}$ being the $m$-th iterate of Algorithm 2 and the $L^{2}({[0,T]})$-norm over the interval $[0,T]$ is as defined in Table 1. The algorithm is always halted when the error $\varepsilon_{m}$ as computed above is below 1%.
@@ 4.1 Kuramoto-Shinomoto-Sakaguchi MV-SDE
Consider the following MV-SDE, which is related to the famous Kuramoto-Shinomoto-Sakaguchi model (see [24, Section 5.3.2]):
@@
We test different choices of algorithm hyperparameters: $r_{0}=1,5,10$, $\rho=0.6,0.7,0.8,0.9$, n=3,4,5,6,7,8n=3,4,5,6{,7,8} and $M=1,10,100,1000,10000$. The benchmark ${\bar{\gamma}}^{\text{MC}}$ in (57), based on the Monte Carlo particle system approximation, is obtained with $N_{0}=10^{6}$ particles. For each combination of hyperparameters, we perform 1000 independent runs of the algorithm, and record the average number of iterations of the SGD algorithm required to achieve the relative error accuracy $\varepsilon_{m}<1\%$.
@@
In Figures 2, 3, 4 we plot the output curves $\mathscr{L}{\bf a}_{m}=(\mathscr{L}{\bf a}_{m}^{1},\mathscr{L}{\bf a}_{m}^{2})$ of the SGD algorithm versus the benchmark curves ${\bar{\gamma}}^{\text{MC}}=({\bar{\gamma}}^{\text{MC}}_{1},{\bar{\gamma}}^{\text{MC}}_{2}),$ for all values of the polynomial degree n=3,4,,8n=3,4,...,8 and for time horizons $T=0.5,1,2,$ respectively. Note that the approximation does not deteriorate as the time horizon $T$ increases. In particular, the slight loss of precision that can be observed towards the right end of the time interval can be explained by the choice of polynomial approximations. For instance, the gap between $(\mathscr{L}{\bf a}_{m})^{1}$ and ${\bar{\gamma}}^{\text{MC}}_{1}$, which can be seen in Figure 2 in the region $t\in[0.4,0.5]$, does not appear in the analogous plot of Figure 3 (or Figure 4) after the same number of steps. This shows that the error does not propagate over time: this is mainly due to the fact that ours is a full history-type minimization algorithm, in that it approximates the minimum in (11) over the set of the continuous curves on $[0,T]$.
@@ 4.2 Polynomial drift MV-SDE
In Figure 5 we plot the output curves $\mathscr{L}{\bf a}_{m}=(\mathscr{L}{\bf a}_{m}^{1},\mathscr{L}{\bf a}_{m}^{2})$ of the SGD algorithm versus the benchmark curves ${\bar{\gamma}}^{\text{MC}}=({\bar{\gamma}}^{\text{MC}}_{1},{\bar{\gamma}}^{\text{MC}}_{2})$ for all time horizons $T=0.1$, $T=0.5$ and $T=1$. In the plots, particularly for $T=0.5$, notice the considerable loss of accuracy in our approximation $\mathscr{L}{\bf a}_{m}^{2}$ of $\bar{\gamma}_{2}(t)=\mathbb{E}[X_{t}]$ near the right boundary of the interval (similar effects were visible in the plots relative to the Kuramoto model). After running additional tests, we do not believe that this effect is due to the polynomial projection, similar to Runge’s type phenomena in polynomial interpolation over compact sets, but we claim that it is rather due to the norm used to define the cost function $F^{2}$ (cf. (12)), in combination with the chosen threshold level we set for $\varepsilon_{m}$ in (58). Indeed, lowering this threshold from to $10^{-2}$ to $10^{-3}$ reduces considerably the error, also near the boundaries of the interval. However, one could look for strategies to distribute the error in a more uniform way across the interval, and obtain a better qualitative result while keeping the same error threshold. One possibility of improvement is to first employ the alternative objective function presented in Remark 2.14, i.e., using a weight function to force the SGD algorithm to be more accurate towards the end of the interval. One can also use a similar weight function to compute a weighted $L^{2}({[0,T]})$ norm for error $\varepsilon_{m}$ in the stopping criterion. Another possibility is to experiment with different choices of basis functions listed in Example 2.16 and Example 2.17 and choose the one with best error performance.
@@ 4.3 Convolution-type MV-SDE
$\varphi_{k}(x)=\overline{H}_{k}(x)\mathrm{e}^{-x^{2}/2},$ $\alpha_{k}(x)=\int_{\mathbb{R}}b(x,y)\varphi_{k}(y)dy=\pi^{1/4}\Bigl(\frac{1}{2}\Bigr)^{k/2}\frac{x^{k}}{\sqrt{k!}}\mathrm{e}^{-x^{2}/4}.$ (67)
@@
In Figure 6 we plot the approximate density $\tilde{w}^{(K),\text{MC}}_{T}(x)$ for $K=3,5,10,20$, with $x$ ranging on the interval [3,4][-3,4] . From this we conclude that the value $K=10$ is high enough in order for $\tilde{w}^{(K)}$ to accurately approximate ${w}_{T}$ over the interval [3,4][-3,4]. In Figure 7 we test the accuracy of the SGD algorithm by plotting $\tilde{w}^{(K),\text{SGD}}_{T}(x)$ versus $\tilde{w}^{(K),\text{MC}}_{T}(x)$, with $K=10$, over the same interval. The parameters of the SGD algorithm were chosen as $n=3$, $M=100$, $r_{0}=5$, $\rho=0.9$, and the algorithm halted after $m=172$ iterations ($35$ seconds of computation time) when the relative error reached $\varepsilon_{m}<1\%$.
Figure 6: Monte Carlo method densities $\tilde{w}^{(K),\text{MC}}_{T}(x)$ over the interval [3,4][-3,4], for $T=1$ and $K=3,5,10,20$. The benchmark vector ${\bar{\gamma}}^{\text{MC}}(T)$ was computed with $N_{0}=10^{7}$ particles. Here $X_{0}\sim\mathcal{N}_{(0,1)}$ and $\sigma=0.1$. In the model (63) we had $X_{0}\sim\mathcal{N}_{(0,1)}$ and $\sigma=0.1$.
Figure 7: Comparison of $\tilde{w}^{(K),\text{SGD}}_{T}(x)$ and $\tilde{w}^{(K),\text{MC}}_{T}(x)$ over the interval [3,4][-3,4], for $T=1$ and $K=10$. The benchmark vector ${\bar{\gamma}}^{\text{MC}}(T)$ was computed with $N_{0}=10^{7}$ particles. The SGD output $\mathscr{L}{\bf a}_{m}(T)$ was obtained after $m=172$ iterations ($35$ seconds of computation time), with parameters $n=3$, $M=100$, $r_{0}=5$, $\rho=0.9$. In the model (63) we had $X_{0}\sim\mathcal{N}_{(0,1)}$ and $\sigma=0.1$.
@@ Appendix A SDE stability estimates
Furthermore, for a given $\kappa>0$ and for any $\rho$ and $\hat{\rho}$ with $\|\rho\|_{\mathcal{S}^{\infty}_{[0,T]}},\|\hat{\rho}\|_{\mathcal{S}^{\infty}_{[0,T]}}\leq\kappa$ we have:
@@
    $\lVert\mathcal{Y}^{\rho}-\mathcal{Y}^{\hat{\rho}}\rVert_{\mathcal{S}^{2p}_{[0,T]}}\leq C_{3}\,\|\rho-\hat{\rho}\|_{\mathbb{L}^{2p}_{[0,T]}}\leq C_{3}\,T^{\frac{1}{2p}}\|\rho-\hat{\rho}\|_{\mathcal{S}^{\infty}_{[0,T]}},$ (76)
@@
The well-posedness of $\mathcal{Y}^{\rho}$ and part (a) of the result are particular cases of [35, Theorem 2.5.7 (p.77)] and [35, Theorem 2.5.10 (p.85)], respectively.
Parts (b) and (c) can be readily obtained by combining (73) with [35, Theorem 2.5.9 (p.83)]. Precisely, from [35, Theorem 2.5.9], there exists $C>0$, only dependent on $\kappa,p,T,\Lambda$ and $K$, such that
@@ B.1 Proof of the time-regularity of $\bar{\gamma}$
Let $X$ be the unique solution to (2), which exists due to Proposition 2.4. A straightforward application of Itô’s formula shows that its time-marginal laws $\mu_{t}(\text{d}x)$ solve the Fokker-Plank equation in distributional form over $0,T]\times\mathbb{R}^{d}$, which is
@@ B.2 Auxiliary stability and moment estimates results
$+\big|\nabla{{\bf h}_{j}}\big(\mathscr{L}a(t)\big)\beta(t,Z^{a}_{t})-\nabla{{\bf h}_{j}}\big(\mathscr{L}a^{\prime}(t)\big)\beta(t,Z^{a^{\prime}}_{t})\big|\Big)\hskip 0.0pt$ (96)
@@
Since ${\bf h},\alpha(t,\cdot),\beta(t,\cdot)$ and their derivatives up to order $2$ are bounded, uniformly w.r.t. $t\in[0,T]$, [36, Theorem 2.3.1 (p.218)] yields
$\partial_{a}Z^{a}_{t}=Y^{a}_{t},\quad a\in\mathbb{R},\quad\text{a.s.},$ (102)
@@
[…]
@@ B.3 Proof of the main auxiliary propositions
[…]
$)\displaystyle\mathbb{E}\big[\big|\varphi\big(Z^{a}_{t}(\bar{\xi},\bar{W})\big)-{\bf h}\big((\mathscr{L}a)(t)\big)\big|^{2}\big]\times\mathbb{E}\big[\big|\varphi^{\prime}\big(Z^{a}_{t}(\tilde{\xi},\tilde{W})\big)Y^{a}_{t}(\tilde{\xi},\tilde{W}$$\big)$- $\partial_{a}{\bf h}\big((\mathscr{L}a)(t)\big)\big|^{2}\big]$
@@
$\hskip-8.0pt\mathbb{E}_{\bar{\mathbb{P}}}\big[{\bf v}_{m+1}-\nabla_{a}G({\bf a}_{m})|\bar{\mathcal{F}}_{m}\big]$ (114)
$\hskip-8.0pt=\mathbb{E}_{\bar{\mathbb{P}}}\Big[v({\bf a}_{m};\xi_{m+1},W_{m+1};\tilde{\xi}_{m+1},\tilde{W}_{m+1})-\mathbb{E}_{\bar{\mathbb{P}}}\big[v(a;\xi_{m+1},W_{m+1};\tilde{\xi}_{m+1},\tilde{W}_{m+1})\big]_{a={\bf a}_{m}}\Big|\bar{\mathcal{F}}_{m}\Big]=0,$ (115)